OpenStudy (anonymous):

the half life of radium-226 is 1590. after how many years will only 50 mg remain

OpenStudy (campbell_st):

do you have a growth or decay model for radium...?

OpenStudy (anonymous):

the first part of the question is "if a sample has a mass of 150 mg find the mass that will remain after 1000 years" i did that part and got 96.9 mg remain

OpenStudy (campbell_st):

ok... so you have the model... I don't know what it is... but I'll use exponential just for the purpose of demonstrating so you have 50mg remaining and k is the decay constant \[50 = 150 e^{-kt}\] then divide both sides by 150 \[\frac{1}{3} = e^{-kt}\] now take the base e log of both sides \[\log_{e}(\frac{1}{3}) = -kt\] divide both sides of the equation by -k so the time take for 150mg or radium to decay to 50mg is \[t = \frac{\log_{e}(\frac{1}{3})}{-k}\] hope it makes sense....

OpenStudy (anonymous):

thanks it helped a lot i got it